OpenStudy (anonymous):

A $3,300.00 principal earns 4% interest, compounded annually. After 3 years, what is the balance in the account?

OpenStudy (anonymous):

$3,712.05 $211,200.00 $3,696.00 $10,296.00

OpenStudy (destinymasha):

Future Value = Principal(1 + i)^n FV = 3300(1.04^3) = $3,712.05